# Panoramic Distortion-Aware Tokenization for Person Detection and Localization in Overhead Fisheye Images

### v3

@@ 4 Experiments
Seidel et al[Seidel_VISAPP_2019] 21.8 59.8 07.6 32.8 28.5 02.3 23.2 19.5 20.2 58.2 07.1 32.2 28.3 03.9 22.4 18.9
Li et al[Li_AVSS_2019] 28.5 63.3 20.1 46.8 24.2 01.3 33.8 29.3 27.2 65.2 21.3 47.2 24.8 01.3 31.8 27.6
Tamura et al[Tamura_WACV_2019] 34.8 72.1 27.7 51.7 38.8 08.7 38.8 32.9 34.2 72.8 28.7 53.7 37.3 06.5 39.5 33.2
RAPiD [Duan_CVPRW_2020] 40.3 77.9 34.8 55.3 41.9 09.2 44.7 37.6 39.2 77.9 35.4 54.8 40.1 07.9 44.2 37.3
OARPD [Qiao_MTA_2024] 45.5 78.8 46.3 - - - - - 45.8 79.4 45.8 - - - - -
Yang et al[Yang_ICCV_2023] 47.4 82.6 48.4 64.1 54.3 14.1 50.8 45.7 46.2 81.1 47.3 66.1 53.5 12.6 49.3 44.9
@@
Seidel et al[Seidel_VISAPP_2019] VISAPP’19 20.9 50.6 10.2 80.6 39.5 53.0 16.1 39.4 09.0 70.9 38.6 50.0
Li et al[Li_AVSS_2019] AVSS’19 34.2 75.7 28.6 86.3 65.4 74.4 25.2 69.9 30.2 81.4 64.5 72.0
Tamura et al[Tamura_WACV_2019] 1 WACV’19 28.8 61.6 23.0 73.1 65.8 69.3 12.8 27.2 10.9 53.0 34.3 41.6
RAPiD [Duan_CVPRW_2020] 1 CVPRW’20 38.3 78.6 30.8 91.8 80.3 85.7 25.0 54.0 17.4 83.7 57.6 68.2
OARPD [Qiao_MTA_2024] 1 MTA’24 11.4 28.9 06.6 80.5 31.3 45.1 04.1 10.0 02.2 28.8 19.0 22.9
Yang et al[Yang_ICCV_2023] 1 ICCV’23 39.7 78.3 37.1 82.9 72.6 77.4 20.5 45.3 16.3 66.8 43.0 52.3
@@
To demonstrate the validity and effectiveness of our approach, we conducted extensive experiments using three large-scale datasets [Duan_CVPRW_2020, Tezcan_WACV_2022, Yang_ICCV_2023], which are composed of overhead fisheye images. We used $\times1024$ input resolution by following the settings given in [Duan_CVPRW_2020, Li_AVSS_2019, Seidel_VISAPP_2019, Tamura_WACV_2019, Yang_ICCV_2023]. For evaluation, we converted the trapezoid bounding boxes into rectangles because no other methods listed in Table 1 produced trapezoid outputs.
@@ 4.2 Person localization
For person localization, we adopt the geometry-based method proposed by Yang et al[Yang_ICCV_2023]. This method estimates person locations by calculating the intersection point between the incident ray from the camera and the ground plane, where the incident ray is determined from the center of the lower edge of a detected person-bounding box.
@@ 4.3 Experimental results
We used the person detection and localization results reported by Yang et al[Yang_ICCV_2023] in the methods of Seidel et al[Seidel_VISAPP_2019], Li et al[Li_AVSS_2019], Tamura et al[Tamura_WACV_2019], RAPiD [Duan_CVPRW_2020], and Yang et al[Yang_ICCV_2023]. We also used the OARPD [Qiao_MTA_2024] results reported in [Qiao_MTA_2024]. Unless otherwise specified, we refer to these results throughout. Because of the lack of previous performance reports, we implemented methods for comparison of Tamura et al[Tamura_WACV_2019] and OARPD [Qiao_MTA_2024], based on the corresponding papers using PyTorch [Paszke_NIPS_2019]. In addition, the official codes from Yang et al[Yang_ICCV_2023] and RAPiD [Duan_CVPRW_2020] were used for evaluation.
@@ 4.3.1 Person detection on LOAF
Metrics and protocols. To demonstrate the validity and the effectiveness of the proposed method on LOAF, we used the mean average precision (mAP) following the COCO evaluation [Lin_ECCV_2014]. We used the subscripts $n$, $m$, and $f$ to denote the near (0–10 m), mid-range (10–20 m), and distant (over 20 m) person positions, respectively, e.g., APn. LOAF also provides seen and unseen scenes for test splits using the subscripts seenseen and unseenunseen, respectively. We used the official splits of the training, validation (val), and test sets.
Results. Our method achieved the highest mAP among the methods listed in Table 2 when used on both the validation and test sets. The mAP obtained using our method was substantially higher than that obtained using Yang et al.’s method [Yang_ICCV_2023] by values of 5.0 and 8.3 on the LOAF validation and test sets, respectively. In addition, the APf value obtained using our method on the test set was notably larger than that of Yang et al.’s method [Yang_ICCV_2023] by 32.0, where 67% of the bounding boxes were labeled as $f$: over 20 m. Our method also achieved an APseen of 56.2 and an APunseen of 53.6 on the test set, thus outperforming the other methods. The substantial margin between our method and Yang et al.’s method [Yang_ICCV_2023] on the unseen test set of 8.7 demonstrates the scene robustness.
@@ 4.3.2 Person detection on CEPDOF
Seidel et al[Seidel_VISAPP_2019] VISAPP’19 1.298 0.561 1.332 3.109 1.206 1.382 1.321 0.706 1.309 3.482 1.306 1.386
Li et al[Li_AVSS_2019] AVSS’19 0.898 0.502 0.871 2.650 0.832 0.962 0.913 0.543 0.884 2.780 0.904 0.998
Tamura et al[Tamura_WACV_2019] WACV’19 0.755 0.429 0.736 1.862 0.709 0.821 0.778 0.471 0.826 2.160 0.724 0.836
RAPiD [Duan_CVPRW_2020] CVPRW’20 0.674 0.426 0.623 1.403 0.625 0.757 0.682 0.461 0.664 1.445 0.638 0.776
OARPD [Qiao_MTA_2024] MTA’24 - - - - - - - - - - - -
Yang et al[Yang_ICCV_2023] ICCV’23 0.368 0.144 0.363 0.775 0.318 0.402 0.374 0.148 0.375 0.819 0.326 0.398
@@
Metrics and protocols. CEPDOF [Duan_CVPRW_2020] provides bounding box annotations without person locations in the world coordinates. Therefore, we evaluated the detection performance using the AP. Similar to prior works [Duan_CVPRW_2020, Tezcan_WACV_2022, Yang_ICCV_2023], we also used precision, recall, and the F1-score. Following the procedure of [Yang_ICCV_2023], we trained the method of Tamura et al[Tamura_WACV_2019], along with RAPiD [Duan_CVPRW_2020], OARPD [Qiao_MTA_2024], Yang et al[Yang_ICCV_2023], and our method, using HABBOF [Li_AVSS_2019] and WEPDTOF [Tezcan_WACV_2022] after pretraining these methods on COCO [Lin_ECCV_2014], although MW-R [Duan_CVPRW_2020] was not used for the training because it is not publicly available.
Results. Our method achieved the highest mAP, AP50, and AP75 values on CEPDOF among the methods listed in Table 3. The mAP, AP50, and AP75 values achieved using our method are higher than those obtained using Yang et al.’s method [Yang_ICCV_2023] by values of 3.1, 7.3, and 1.6, respectively. Additionally, the F1-score of our method was greater than that of Yang et al.’s method [Yang_ICCV_2023] by 9.1.
@@ 4.3.3 Person detection on WEPDTOF
Metrics and protocols. We used the evaluation metrics of CEPDOF, as described in Section 4.3.2. We trained the methods of Tamura et al[Tamura_WACV_2019], RAPiD [Duan_CVPRW_2020], OARPD [Qiao_MTA_2024], Yang et al[Yang_ICCV_2023], and our method, using HABBOF and CEPDOF after they were pretrained on COCO following [Yang_ICCV_2023], although MW-R was again not used for the training because it is not publicly available.
Results. Similar to the CEPDOF results, our method achieved the highest mAP, AP50, and AP75 results on WEPDTOF among the methods listed in Table 3. The mAP values achieved on CEPDOF using both the conventional methods and our method were larger than those obtained on WEPDTOF because WEPDTOF has more scenes; i.e., WEPDTOF consists of 16 videos, and CEPDOF consists of 8 videos. Comparison between the CEPDOF and WEPDTOF results shows that the methods of Tamura et al[Tamura_WACV_2019], RAPiD [Duan_CVPRW_2020], OARPD [Qiao_MTA_2024], and Yang et al[Yang_ICCV_2023] apparently require diverse images because the network is trained on fisheye images that include rotated persons.
@@ 4.3.4 Diagnostic study
We conducted a diagnostic study on the LOAF test set. Table 4 shows the ablation study results. The baseline represents DAB-DETR [Liu_ICLR_2022] with the SG-Former-S [Ren_ICCV_2023] for the backbone when using fisheye input images. The second row demonstrates the effectiveness of the panoramic conversion. The panoramic images were suitable for DAB-DETR because they removed the person rotations. After the panoramic conversion was adopted, the performance increased by an mAP of 9.6. We also validated the effectiveness of the PDAT process. The third row indicates that our method using PDAT provided substantially improved person detection results when compared with our method without PDAT by an mAP of 4.7. Therefore, the PDAT alleviated the problems caused by the variations in person size dramatically. Similar to the mAP, the other metrics were improved notably.
@@ 4.3.5 Person localization
Metrics and protocols. To validate the person localization accuracy, we compared our method with the conventional methods. We evaluated the individual person positions from the bounding boxes using Yang et al.’s approach [Yang_ICCV_2023], as described in Section 4.2. We used the position error (PE) in meters as the horizontal distance between the ground-truth person position and the person position estimated from the bounding boxes. We also used the same subscripts used in Section 4.3.1, e.g., PEn. These estimated bounding boxes were the same as the person detection results on LOAF from Section 4.3.1.
Results. Our method achieved the lowest mean PE (mPE), as indicated in Table 6. The mPE for our method is smaller than that of Yang et al.’s method [Yang_ICCV_2023] on the LOAF test set by 0.097 m. Similar to the person detection performance, our method realized a noteworthy 0.124-m gain in terms of the PEunseen when compared with Yang et al.’s method [Yang_ICCV_2023]. Our method outperformed the other methods on PEn, PEm, and PEf because of the accuracy of bounding box detection.
@@ 5 Conclusion
Limitations. We assume that the fisheye images are stereographic projections. However, this assumption may cause a mismatch with the actual distortion. To address this problem, we will extend our method using deep single image camera calibration techniques [Dal_ECCV_2024, Wakai_ECCV_2022] that estimate the lens distortion. Another promising direction for future work is to use videos for the input. In this paper, we have focused on person detection and localization from an image.

### v4

@@ 4 Experiments
Seidel et al. [48] 21.8 59.8 07.6 32.8 28.5 02.3 23.2 19.5 20.2 58.2 07.1 32.2 28.3 03.9 22.4 18.9
Li et al. [26] 28.5 63.3 20.1 46.8 24.2 01.3 33.8 29.3 27.2 65.2 21.3 47.2 24.8 01.3 31.8 27.6
Tamura et al. [55] 34.8 72.1 27.7 51.7 38.8 08.7 38.8 32.9 34.2 72.8 28.7 53.7 37.3 06.5 39.5 33.2
RAPiD [11] 40.3 77.9 34.8 55.3 41.9 09.2 44.7 37.6 39.2 77.9 35.4 54.8 40.1 07.9 44.2 37.3
OARPD [39] 45.5 78.8 46.3 - - - - - 45.8 79.4 45.8 - - - - -
Yang et al. [73] 47.4 82.6 48.4 64.1 54.3 14.1 50.8 45.7 46.2 81.1 47.3 66.1 53.5 12.6 49.3 44.9
@@
Seidel et al. [48] VISAPP’19 20.9 50.6 10.2 80.6 39.5 53.0 16.1 39.4 09.0 70.9 38.6 50.0
Li et al. [26] AVSS’19 34.2 75.7 28.6 86.3 65.4 74.4 25.2 69.9 30.2 81.4 64.5 72.0
Tamura et al. [55] 1 WACV’19 28.8 61.6 23.0 73.1 65.8 69.3 12.8 27.2 10.9 53.0 34.3 41.6
RAPiD [11] 1 CVPRW’20 38.3 78.6 30.8 91.8 80.3 85.7 25.0 54.0 17.4 83.7 57.6 68.2
OARPD [39] 1 MTA’24 11.4 28.9 06.6 80.5 31.3 45.1 04.1 10.0 02.2 28.8 19.0 22.9
Yang et al. [73] 1 ICCV’23 39.7 78.3 37.1 82.9 72.6 77.4 20.5 45.3 16.3 66.8 43.0 52.3
@@
To demonstrate the validity and effectiveness of our approach, we conducted extensive experiments using three large-scale datasets [11, 56, 73], which are composed of overhead fisheye images. We used $\times1024$ input resolution by following the settings given in [11, 26, 48, 55, 73]. For evaluation, we converted the trapezoid bounding boxes into rectangles because no other methods listed in Table 1 produced trapezoid outputs. The experimental details are given in the supplementary material.
@@ 4.2 Person localization
For person localization, we adopt the geometry-based method proposed by Yang et al. [73]. This method estimates person locations by calculating the intersection point between the incident ray from the camera and the ground plane, where the incident ray is determined from the center of the lower edge of a detected person-bounding box.
@@ 4.3 Experimental results
We used the person detection and localization results reported by Yang et al. [73] in the methods of Seidel et al. [48], Li et al. [26], Tamura et al. [55], RAPiD [11], and Yang et al. [73]. We also used the OARPD [39] results reported in [39]. Unless otherwise specified, we refer to these results throughout. Because of the lack of previous performance reports, we implemented methods for comparison of Tamura et al. [55] and OARPD [39], based on the corresponding papers using PyTorch [37]. In addition, the official codes from Yang et al. [73] and RAPiD [11] were used for evaluation. A qualitative evaluation is given in the supplementary material.
@@ 4.3.1 Person detection on LOAF
Metrics and protocols. To demonstrate the validity and the effectiveness of the proposed method on LOAF, we used the mean average precision (mAP) following the COCO evaluation [29]. We used the subscripts $n$, $m$, and $f$ to denote the near (0–10 m), mid-range (10–20 m), and distant (over 20 m) person positions, respectively, e.g., APn. LOAF also provides seen and unseen scenes for test splits using the subscripts seenseen and unseenunseen, respectively. We used the official splits of the training, validation (val), and test sets.
Results. Our method achieved the highest mAP among the methods listed in Table 2 when used on both the validation and test sets. The mAP obtained using our method was substantially higher than that obtained using Yang et al.’s method [73] by values of 5.0 and 8.3 on the LOAF validation and test sets, respectively. In addition, the APf value obtained using our method on the test set was notably larger than that of Yang et al.’s method [73] by 32.0, where 67% of the bounding boxes were labeled as $f$: over 20 m. Our method also achieved an APseen of 56.2 and an APunseen of 53.6 on the test set, thus outperforming the other methods. The substantial margin between our method and Yang et al.’s method [73] on the unseen test set of 8.7 demonstrates the scene robustness.
@@ 4.3.2 Person detection on CEPDOF
Seidel et al. [48] VISAPP’19 1.298 0.561 1.332 3.109 1.206 1.382 1.321 0.706 1.309 3.482 1.306 1.386
Li et al. [26] AVSS’19 0.898 0.502 0.871 2.650 0.832 0.962 0.913 0.543 0.884 2.780 0.904 0.998
Tamura et al. [55] WACV’19 0.755 0.429 0.736 1.862 0.709 0.821 0.778 0.471 0.826 2.160 0.724 0.836
RAPiD [11] CVPRW’20 0.674 0.426 0.623 1.403 0.625 0.757 0.682 0.461 0.664 1.445 0.638 0.776
OARPD [39] MTA’24 - - - - - - - - - - - -
Yang et al. [73] ICCV’23 0.368 0.144 0.363 0.775 0.318 0.402 0.374 0.148 0.375 0.819 0.326 0.398
@@
Metrics and protocols. CEPDOF [11] provides bounding box annotations without person locations in the world coordinates. Therefore, we evaluated the detection performance using the AP. Similar to prior works [11, 56, 73], we also used precision, recall, and the F1-score. Following the procedure of [73], we trained the method of Tamura et al. [55], along with RAPiD [11], OARPD [39], Yang et al. [73], and our method, using HABBOF [26] and WEPDTOF [56] after pretraining these methods on COCO [29], although MW-R [11] was not used for the training because it is not publicly available.
Results. Our method achieved the highest mAP, AP50, and AP75 values on CEPDOF among the methods listed in Table 3. The mAP, AP50, and AP75 values achieved using our method are higher than those obtained using Yang et al.’s method [73] by values of 3.1, 7.3, and 1.6, respectively. Additionally, the F1-score of our method was greater than that of Yang et al.’s method [73] by 9.1.
@@ 4.3.3 Person detection on WEPDTOF
Metrics and protocols. We used the evaluation metrics of CEPDOF, as described in Section 4.3.2. We trained the methods of Tamura et al. [55], RAPiD [11], OARPD [39], Yang et al. [73], and our method, using HABBOF and CEPDOF after they were pretrained on COCO following [73], although MW-R was again not used for the training because it is not publicly available.
Results. Similar to the CEPDOF results, our method achieved the highest mAP, AP50, and AP75 results on WEPDTOF among the methods listed in Table 3. The mAP values achieved on CEPDOF using both the conventional methods and our method were larger than those obtained on WEPDTOF because WEPDTOF has more scenes; i.e., WEPDTOF consists of 16 videos, and CEPDOF consists of 8 videos. Comparison between the CEPDOF and WEPDTOF results shows that the methods of Tamura et al. [55], RAPiD [11], OARPD [39], and Yang et al. [73] apparently require diverse images because the network is trained on fisheye images that include rotated persons.
@@ 4.3.4 Diagnostic study
We conducted a diagnostic study on the LOAF test set. Table 4 shows the ablation study results. The baseline represents DAB-DETR [31] with the SG-Former-S [44] for the backbone when using fisheye input images. The second row demonstrates the effectiveness of the panoramic conversion. The panoramic images were suitable for DAB-DETR because they removed the person rotations. After the panoramic conversion was adopted, the performance increased by an mAP of 9.6. We also validated the effectiveness of the PDAT process. The third row indicates that our method using PDAT provided substantially improved person detection results when compared with our method without PDAT by an mAP of 4.7. Therefore, the PDAT alleviated the problems caused by the variations in person size dramatically. Similar to the mAP, the other metrics were improved notably.
@@ 4.3.5 Person localization
Metrics and protocols. To validate the person localization accuracy, we compared our method with the conventional methods. We evaluated the individual person positions from the bounding boxes using Yang et al.’s approach [73], as described in Section 4.2. We used the position error (PE) in meters as the horizontal distance between the ground-truth person position and the person position estimated from the bounding boxes. We also used the same subscripts used in Section 4.3.1, e.g., PEn. These estimated bounding boxes were the same as the person detection results on LOAF from Section 4.3.1.
Results. Our method achieved the lowest mean PE (mPE), as indicated in Table 6. The mPE for our method is smaller than that of Yang et al.’s method [73] on the LOAF test set by 0.097 m. Similar to the person detection performance, our method realized a noteworthy 0.124-m gain in terms of the PEunseen when compared with Yang et al.’s method [73]. Our method outperformed the other methods on PEn, PEm, and PEf because of the accuracy of bounding box detection.
@@ 5 Conclusion
Limitations. We assume that the fisheye images are stereographic projections. However, this assumption may cause a mismatch with the actual distortion. To address this problem, we will extend our method using deep single image camera calibration techniques [5, 62] that estimate the lens distortion. Another promising direction for future work is to use videos for the input. In this paper, we have focused on person detection and localization from an image.